# Optimistic Query Routing in Clustering-based Approximate Maximum Inner Product Search

### v3

@@ 2.5. Proofs and justification of assumptions
We start by proving Lemma 3. We then proceed to justify the assumptions of the lemma in the subsequent section.
2.5.1. Proof of Lemma 3
@@
(9) $=\sup_{\lVert v\rVert_{2}=1}\big{\lvert}v^{\top}(R_{\circ}-[Q]_{t}[\Lambda]_{t}[Q]_{t}^{\top})v\big{\rvert}$
(10) $=\sup_{\lVert v\rVert_{2}=1}\big{\lvert}\sum_{i=t+1}^{d}\Lambda_{i}\cdot\langle Q_{i},v\rangle^{2}\big{\rvert}$
(11) $=\max_{l\in[t+1,d]}\lvert\Lambda_{l}\rvert.$
@@
(12) $\frac{\text{err}(\Sigma,\Sigma_{t}^{\textsc{ms}})}{\|D\|_{\infty}}\leq\max_{l\in[t+1,d]}\lvert\Lambda_{l}\rvert.$
@@
(13) $\frac{\text{err}(\Sigma,\Sigma_{t}^{\textsc{ms}})}{\|D\|_{\infty}}$ $\leq\max_{l\in[t+1,d]}|\Lambda_{l}|$
(14) $\leq 1+\Lambda_{t+1}$
(15) $\leq\max_{l\in[t+1,d]}|1+\Lambda_{l}|$
(16) $=\text{err}(\underbrace{I+R_{\circ}}_{D^{-1/2}\Sigma D^{-1/2}},[Q]_{t}[I+\Lambda]_{t}[Q]_{t}^{\top}).$
@@
(17) $\text{err}(D^{-1/2}$ $\Sigma D^{-1/2},[Q]_{t}[I+\Lambda]_{t}[Q]_{t}^{\top})$
(18) $\leq\text{err}(D^{-1/2}\Sigma D^{-1/2},D^{-1/2}\Sigma_{t}^{\textsc{lr}}D^{-1/2})$
(19) $=\sup_{\lVert v\rVert_{2}=1}\big{\lvert}v^{\top}D^{-1/2}(\Sigma-\Sigma_{t}^{\textsc{lr}})D^{-1/2}v\big{\rvert}$
(20) $=\sup_{\lVert v\rVert_{2}=1}\big{\lvert}\sum_{i,j}v_{i}v_{j}(\Sigma-\Sigma_{t}^{\textsc{lr}})_{ij}/\sqrt{D_{ii}D_{jj}}\big{\rvert}$
(21) $\leq\sup_{\lVert v\rVert_{2}=1}\big{\lvert}\sum_{i,j}v_{i}v_{j}(\Sigma-\Sigma_{t}^{\textsc{lr}})_{ij}\frac{1}{(1-\epsilon)\lVert D\rVert_{\infty}}\big{\rvert}$
(22) $\leq\frac{1}{(1-\epsilon)\lVert D\rVert_{\infty}}\sup_{\lVert v\rVert_{2}=1}\big{\lvert}v^{\top}(\Sigma-\Sigma_{t}^{\textsc{lr}})v\big{\rvert}$
(23) $=\frac{1}{(1-\epsilon)\lVert D\rVert_{\infty}}\text{err}(\Sigma,\Sigma_{t}^{\textsc{lr}}),$
@@
Figure 2. Histogram of the $(t+1)$-th eigenvalue. For each dataset, we pick the partitioning and $t$ from Table 1. Plots show that almost all shards for all datasets have $(t+1)$-th eigenvalue bounded away from $1$, except for a few shards for Nq-Ada2.
2.5.2. Assumptions of Lemma 3
@@
While in the worst case, the diagonal of $\Sigma$ can have arbitrarily large entries compared to its smallest entries, in practice this is rarely the case. While we do not explore how to remove this assumption, there are several mechanisms to do so in practice such as applying random rotations or pseudo-random rotations (Ailon and Chazelle, 2009; Woodruff, 2014; Ailon and Liberty, 2013) to the data points in each shard before using Algorithm 1. It is well known (e.g. Lemma 1 in (Ailon and Chazelle, 2009)) that after applying such transforms, the coordinates of the vectors are “roughly equal,” thereby ensuring that the diagonal of the covariance has entries of comparable magnitude. We leave the exploration of removing this assumption to future work.
3. Experimental evaluation
We put our arguments to the test and evaluate Optimist.
3.1. Setup
Datasets: We use the following benchmark datasets in this work:
  • Text2Image: A cross-modal dataset, where data and query points may have different distributions in a shared space (Simhadri et al., 2022). We use a subset consisting of $10$ million $200$-dimensional data points along with a subset of $10{,}000$ test queries;
  • Music: $1$ million $100$-dimensional points (Morozov and Babenko, 2018) with $1{,}000$ queries;
  • DeepImage: Subset of $10$ million $96$-dimensional points from the billion deep image features dataset (Yandex and Lempitsky, 2016) with $10{,}000$ queries;
  • GloVe: 1.21.21{.}21.2 million, $200$-dimensional word embeddings trained on tweets (Pennington et al., 2014) with $10{,}000$ test queries;
  • MsMarco-MiniLM: Ms Marco Passage Retrieval v1 (Nguyen et al., 2016) is a question-answering dataset consisting of 8.88.88{.}88.8 million short passages in English. We use the “dev” set of queries for retrieval, made up of $6{,}980$ questions. We embed individual passages and queries using the all-MiniLM-L6-v2 model [^6] Checkpoint at https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. to form a $384$-dimensional vector collection; and,
  • Nq-Ada2: 2.72.72{.}72.7 million, $1{,}536$-dimensional embeddings of the Natural Questions dataset (Kwiatkowski et al., 2019) with the Ada-002 model. [^7] https://openai.com/index/new-and-improved-embedding-model/
@@ 3.1. Setup
Table 1. Dataset statistics (size $m$ and number of dimensions $d$), along with the number of partitions ($C$) and Optimist’s default rank configuration ($t$) in our main experiments.
Dataset $m$ $d$ $C$ $t$
Text2Image $10$M $200$ $3{,}000$ $4$
Music $1$M $001$,02410241{,}0241 , 024 $2$
DeepImage $10$M $96$ $3{,}000$ $2$
GloVe $1.2$M $200$ $1{,}024$ $4$
MsMarco-MiniLM $8.8$M $843$,00030003{,}0003 , 000 $8$
Nq-Ada2 $2.7$M $1{,}536$ $1{,}600$ $30$
Clustering: For our main results, we partition datasets with Spherical KMeans (Dhillon and Modha, 2001). We include in Appendix A results from similar experiments but where the clustering algorithm is standard KMeans and Gaussian Mixture Model (GMM). We cluster each dataset into $\sqrt{m}$ shards, where $m$ is the number of data points in the dataset; this is summarized in Table 1.
Evaluation: Once a dataset has been partitioned, we fix the partitioning and evaluate all routers on it to facilitate a fair comparison. We evaluate each router $\tau$ as follows. For each test query, we identify the set of shards to probe using $\tau$. We then perform exact search over the selected shards, obtain the top-$k$ points, and compute recall with respect to the ground-truth top-$k$ set. Because the only source of error is the router’s inaccuracy, the measured recall gauges the effectiveness of $\tau$.
Routers: We evaluate the following routers in our experiments:
  • Mean and NormalizedMean: Defined in Equations (2) and (3);
  • Scann $(T)$: Similar to Mean and NormalizedMean, but where routing is determined by inner product between a query and the Scann centroids (c.f., Theorem 4.24.24{.}24.2 in (Guo et al., 2020)). Scann has a single hyperparameter $T$, which we set to $0.5$ after tuning;
  • SubPartition $(t)$: Recall that Optimist stores $t+2$ vectors per partition, where $t$ is the rank in Algorithm 1. We ask if simply partitioning each shard independently into $t+2$ sub-partitions, and recording the sub-partitions’ centroids as the representatives of that shard attains the same routing accuracy as Optimist. At query time, we take the maximum inner product of the query with a shard’s representatives as its score, and sort shards according to this score; and,
  • Optimist $(t,\delta)$: The Optimist router given in Algorithm 1. The parameter $t$ determines the rank of the sketch of the covariance matrix, and $\delta$ determines the degree of optimism. We set $\delta=0.8$ and $t$ to a maximum of $2\%$ of $d$, as summarized in Table 1, but study their effect later. Throughout our discussion, if we do not specify the parameters, it should be understood that Optimist is configured with default $\delta$ and $t$. We write Optimist $(t=d,\cdot)$ to indicate that the full covariance matrix is used.
(a) Nq-Ada2
(b) Music
(c) GloVe
Figure 3. Top-$k$ recall vs. volume of probed data points, on Nq-Ada2, Music, and GloVe. Partitioning is by Spherical KMeans. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$.
(a) Text2Image
(b) MsMarco-MiniLM
(c) DeepImage
Figure 4. Top-$k$ recall vs. volume of probed data points, on Text2Image, MsMarco-MiniLM, and DeepImage. Setup is the same as Figure 3.
3.2. Main results
Figures 3 and 4 plot top-$k$ recall for $k\in\{1,10,100\}$, versus the percentage of data points examined for all datasets. Partitioning is by Spherical KMeans; for standard KMeans and GMM see Figures 9 and 10.
We summarize key observations. First, among baselines, NormalizedMean generally outperforms Mean and Scann, except Music where Mean reaches a higher recall. Second, with very few exceptions, Optimist with the full covariance ($t=d$) does at least as well as NormalizedMean, and often outperforms it significantly. Interestingly, the gap between baselines and Optimist widens as retrieval depth ($k$) increases; a phenomenon that is not surprising.
Finally, while Optimist with a rank-$t$ Masked Sketch shows some degradation in comparison with no sketching (i.e., $t=d$), it still achieves a higher recall than baselines for larger $k$. When $k$ is smaller, SubPartition $(t)$ becomes a strong competitor.
Optimist shines when data points have varying norms. On Music, at $95\%$ top-$100$ recall, Optimist needs to probe $54\%$ fewer data points than NormalizedMean; on average Optimist probes $66{,}6$ 666 data points to reach $95\%$ top-$100$ recall whereas NormalizedMean examines $14{,}463$ points.
Table 2 presents savings on all datasets. On DeepImage, Optimist scans $9\%$ more data points than NormalizedMean to reach $90\%$ recall ($55{,}000$ points vs $60{,}000$), and $10\%$ more to reach $95\%$ ($95{,}000$ vs $105{,}000$). We suspect this is an artifact of the dataset’s construction: each point is represented by the top $96$ principal components of its original features, leading to unusual partition statistics.
Table 2. Comparison of Optimist and NormalizedMean by relative amount of data points saved during search to achieve a fixed recall. A saving of $x\%$ means that Optimist searches $x\%$ fewer points than NormalizedMean.
Recall Nq-Ada2 GloVe MsMarco-MiniLM Music Text2Image
$90\%$ $18\%$ $11\%$ $22\%$ $38\%$ $23\%$
$95\%$ $20\%$ $5.5\%$ $7.7\%$ $54\%$ $22\%$
3.2.1. Latency
As Table 2 shows, Optimist reaches $95\%$ recall by processing up to $54\%$ fewer points. We now study the impact of the improved routing accuracy on query latency, which consists of routing, fetching the chosen shards from storage, and scoring.
Concretely, we isolate routing and measure the elapsed time between when a router receives a query and when it returns a list of shards. This is the “routing” latency. We then measure the time it takes to fetch the chosen shards from storage; the “fetch” latency. Finally, we measure the latency of the scoring phase. We note that, for latency experiments only, we compress data points and compute inner products using PQ with $4$-dimensional codebooks.
We run experiments on an AWS c5.xlarge machine with $4$ vCPUs, $8$GB of main memory. Baseline bandwidth of the SSD attached to this machine is $1{,}150$Mbps. [^9] See official benchmark from Amazon at https://docs.aws.amazon.com/AWSEC2/latest/UserGuide/ebs-optimized.html Finally, using $4$ threads, transferring $4$MB of data from blob storage (hosted on Amazon S3) to main memory has P50 latency of $45$ milliseconds (ms). [^10] See https://github.com/dvassallo/s3-benchmark for an independent benchmark.
Refer to caption Refer to caption Refer to caption
(a) $1$ thread
(b) $2$ threads
(c) $4$ threads
Figure 5. Mean per-query latency (milliseconds) to reach $95\%$ recall, using different number of threads (columns), when data is stored on SSD (top row) and blob storage (bottom row). For each dataset, we plot the latency breakdown for NormalizedMean (top bar) and Optimist (bottom bar), and report relative gains (negative values indicate improvements, whereas positive values indicate losses) achieved by Optimist with respect to NormalizedMean. All shards are stored as PQ-compressed blobs using $4$-dimensional codebooks—a common configuration.
Figure 5 reports the routing, fetch, and scoring latency for Optimist and NormalizedMean routers. As the figure shows, Optimist’s gains are more pronounced when shards are stored on blob storage, but even on SSD the gains are substantial.
We note that, hardware plays an outsize role in latency improvements. Depending on the specifications, we may have a slower or faster interface with blob or SSD storage; a more limited bandwidth would boost Optimist’s standing in latency comparisons. On the other hand, if storage bandwidth is abundant and number of threads very limited, Optimist’s advantage would become less significant.
There are other factors that may affect the overall latency comparison. These include: target recall level; type of compression; and algorithm used for scoring. In the above experiments, we put Optimist at a disadvantage by incorporating heavy compression and fast score computation; Optimist fares even better if data remains uncompressed (as Optimist needs to transfer a smaller amount of data) or scoring is slower (as Optimist must score fewer points).
3.2.2. Router size
As we noted, a router such as NormalizedMean that uses a point estimate to score shards needs to store only a single $d$-dimensional vector per shard. Optimist, on the other hand, requires $(t+2)$ vectors in $\mathbb{R}^{d}$ per shard to make its routing decision. Here, we present concrete numbers for completeness, and compare the size of the two routers on the datasets used in our experiments.
On most datasets, the difference in size between NormalizedMean and Optimist is negligible. On GloVe, for example, NormalizedMean takes up $0.8$MB of space while Optimist $(t=4)$ consumes $4.7$MB in total. Statistics on other datasets are as follows: on MsMarco-MiniLM it is $4.4$MB for NormalizedMean versus $44$MB for Optimist $(t=8)$; on Music, $0.4$MB versus $1.6$MB ($t=2$); on Text2Image, $2.3$MB versus $13.7$MB ($t=4$); on DeepImage, $1.1$MB versus $4.4$MB ($t=2$). The difference is larger, but still mild, on Nq-Ada2: $9.4$MB versus $300$MB ($t=30$).
3.3. Effect of parameters on Optimist
@@
Figure 6. Top-$100$ recall vs. volume of probed data as we change the rank of the covariance sketch ($t$). $\textsc{Optimist}(0,\cdot)$ denotes a sketch that only retains the diagonals. Partitioning is with Spherical KMeans.
@@
Figure 7. Top-$100$ recall vs. volume of probed data, comparing a range of values of $\delta$. As $\delta\rightarrow 1$, Optimist becomes more optimistic. Partitioning is with Spherical KMeans.
@@
3.4. Maximum inner product prediction
@@
Figure 8. Mean prediction error $\mathcal{E}_{\ell}(\tau,\cdot)$, defined in Equation (24), versus $\ell$ (expressed as percent of total number of shards), for various routers and datasets.
We claimed that Optimist is statistically principled. We expect then, that it should give more accurate estimates of the maximum inner product for any given query-partition pair. We examine that claim in this section and quantify the prediction error for all routers.
Fix a dataset, whose $C$ partitions are denoted by $\mathcal{P}_{i}$ for $i\in[C]$, along with a router $\tau$ and query $q$. We write $\tau_{i}$ to denote the score computed by $\tau$ for partition $\mathcal{P}_{i}$ and $q$; for example, NormalizedMean computes $\langle q,\mu_{i}/\lVert\mu_{i}\rVert_{2}\rangle$ and Optimist $\theta_{i}$ per Equation (6).
Note that, the $\tau_{i}$’s induce an ordering among partitions. We denote this permutation by $\pi$, so that $\tau_{\pi_{i}}\geq\tau_{\pi_{i+1}}$. We quantify the prediction error for $\ell\in[C]$ as:
(24) $\mathcal{E}_{\ell}(\tau,q)=\frac{1}{\ell}\sum_{i=1}^{\ell}\left|\frac{\tau_{\pi_{i}}}{\max_{u\in\mathcal{P}_{\pi_{i}}}\langle q,u\rangle}-1\right|.$
This error is 00 when scores produced by the router perfectly match the maximum inner product. The role of $\ell$ is to allow us to factor in the rank of partitions in our characterization of the prediction error. In other words, we can measure the error only for the top $\ell$ shards according to $\tau$. In this way, if we decide that it is not imperative for a router to accurately predict the maximum inner product in low-ranking shards, we can reflect that choice in our calculation.
We measure Equation (24) on all datasets partitioned by Spherical KMeans, and all routers considered in this work. The results are shown in Figure 8, where for each choice of $\ell$, we plot $\mathbb{E}_{q}[\mathcal{E}_{\ell}(\tau,q)]$ for test queries.
[…]
4. Concluding remarks
We studied clustering-based maximum inner product search, where data points are clustered into non-overlapping shards to form an index. At query time, two independent components work together to process a query: a router computes a score for each shard with respect to the query and identifies the subset of top shards to probe; and a subsequent scoring stage that computes (approximate) scores for points in the chosen shards. In this general setup, we considered a storage-backed system where shards do not rest in the main memory, but are stored on external storage and must be fetched into the main memory when a router identifies them. This is an important paradigm for vector search at scale.
Within this framework, we focused on routing. We motivated our work by an unusual but unsurprising observation: NormalizedMean computes a score for each shard that over-estimates the maximum inner product between an arbitrary query and the points in that shard. Interestingly, the extent of over-estimation correlates with the variance of the shard: The more spread-out the points are, the more optimistic the NormalizedMean router becomes.
This behavior is unpredictable, however. It is also hard to explain why it leads to more accurate routing when data points have non-constant norms. But we found the insight that variance should play a role in routing to be worth investigating. We formalized that insight and presented a principled optimistic algorithm whose score for a shard is a more accurate estimate of the maximum inner product—as confirmed by evidence in Figure 8.
As Figure 5 shows, our algorithm, Optimist, is more suitable for settings where shards rest on some external, high-latency storage. That is because, in such settings, spending more computation on routing is acceptable if we end up transferring a smaller amount of data from storage to main memory. It is particularly attractive when the storage medium is a cheap, high-latency blob storage.
Optimist has other implications beyond latency. Needing to transfer a smaller volume of data to memory means that each query requires a smaller chunk of the in-memory cache. As a result, cache can be shared among more queries, leading to a higher throughput in systems where memory availability is the main bottleneck.
Our research takes a first step in exploring unsupervised routing for clustering-based MIPS, and identifies a trade-off space that has not been explored previously. We have more to do, however. We leave to future work an exploration of a more compact sketch of the covariance matrix; and, an efficient realization of our general solution outlined in Section 2.3.
Appendix A Additional experimental results
(a) Nq-Ada2
(b) Music
(c) GloVe
Figure 9. Top-$k$ recall vs. volume of probed data. Partitioning is with Standard KMeans. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$.
(d) Text2Image
(e) MsMarco-MiniLM
(f) DeepImage
Figure 9. Top-$k$ recall vs. volume of probed data. Partitioning is with Standard KMeans. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$.
(a) Music
(b) GloVe
(c) Text2Image
Figure 10. Top-$k$ recall vs. volume of probed data. Partitioning is with Gaussian Mixture Model. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$. We note that, due to the dimensionality of Nq-Ada2, we were unable to complete GMM clustering on this particular dataset.
(d) MsMarco-MiniLM
(e) DeepImage
Figure 10. Top-$k$ recall vs. volume of probed data. Partitioning is with Gaussian Mixture Model. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$. We note that, due to the dimensionality of Nq-Ada2, we were unable to complete GMM clustering on this particular dataset.

### v4

@@ Appendix C Proof of Lemma 2 and justification of its assumptions
We start by proving Lemma 2, then justify the assumptions of the lemma.
C.1 Proof of Lemma 2
@@
$=\sup_{\lVert v\rVert_{2}=1}\big\lvert v^{\top}(R_{\circ}-[Q]_{t}[\Lambda]_{t}[Q]_{t}^{\top})v\big\rvert$ (10)
$=\sup_{\lVert v\rVert_{2}=1}\big\lvert\sum_{i=t+1}^{d}\Lambda_{i}\cdot\langle Q_{i},v\rangle^{2}\big\rvert$ (11)
$=\max_{l\in[t+1,d]}\lvert\Lambda_{l}\rvert$. (12)
@@
$\frac{\text{err}(\Sigma,\Sigma_{t}^{\textsc{ms}})}{\|D\|_{\infty}}\leq\max_{l\in[t+1,d]}\lvert\Lambda_{l}\rvert$. (13)
@@
$\frac{\text{err}(\Sigma,\Sigma_{t}^{\textsc{ms}})}{\|D\|_{\infty}}$ $\leq\max_{l\in[t+1,d]}|\Lambda_{l}|\leq 1+\Lambda_{t+1}\leq\max_{l\in[t+1,d]}|1+\Lambda_{l}|$ (14)
$=\text{err}(\underbrace{I+R_{\circ}}_{D^{-1/2}\Sigma D^{-1/2}},[Q]_{t}[I+\Lambda]_{t}[Q]_{t}^{\top}).$ (15)
@@
$\text{err}(D^{-1/2}$ $\Sigma D^{-1/2},[Q]_{t}[I+\Lambda]_{t}[Q]_{t}^{\top})$ (16)
$\leq\text{err}(D^{-1/2}\Sigma D^{-1/2},D^{-1/2}\Sigma_{t}^{\textsc{lr}}D^{-1/2})$ (17)
$=\sup_{\lVert v\rVert_{2}=1}\big\lvert v^{\top}D^{-1/2}(\Sigma-\Sigma_{t}^{\textsc{lr}})D^{-1/2}v\big\rvert$ (18)
$=\sup_{\lVert v\rVert_{2}=1}\big\lvert\sum_{i,j}v_{i}v_{j}(\Sigma-\Sigma_{t}^{\textsc{lr}})_{ij}/\sqrt{D_{ii}D_{jj}}\big\rvert$ (19)
$\leq\sup_{\lVert v\rVert_{2}=1}\big\lvert\sum_{i,j}v_{i}v_{j}(\Sigma-\Sigma_{t}^{\textsc{lr}})_{ij}\frac{1}{(1-\epsilon)\lVert D\rVert_{\infty}}\big\rvert$ (20)
$\leq\frac{1}{(1-\epsilon)\lVert D\rVert_{\infty}}\sup_{\lVert v\rVert_{2}=1}\big\lvert v^{\top}(\Sigma-\Sigma_{t}^{\textsc{lr}})v\big\rvert$ (21)
$=\frac{1}{(1-\epsilon)\lVert D\rVert_{\infty}}\text{err}(\Sigma,\Sigma_{t}^{\textsc{lr}}),$ (22)
@@
Figure 5: Histogram of the $(t+1)$-th eigenvalue. For each dataset, we pick the partitioning and $t$ from Table 2. Plots show that almost all shards for all datasets have $(t+1)$-th eigenvalue bounded away from 11, except for a few shards for Nq-Ada2.
C.2 Assumptions of Lemma 2
@@
While in the worst case, the diagonal of $\Sigma$ can have arbitrarily large entries compared to its smallest entries, in practice this is rarely the case. While we do not explore how to remove this assumption, there are several mechanisms to do so in practice such as applying random rotations or pseudo-random rotations Ailon and Chazelle [2009], Woodruff [2014], Ailon and Liberty [2013] to the data points in each shard before using Algorithm 1. It is well known (e.g. Lemma 1 in Ailon and Chazelle [2009]) that after applying such transforms, the coordinates of the vectors are “roughly equal,” thereby ensuring that the diagonal of the covariance has entries of comparable magnitude. We leave the exploration of removing this assumption to future work.
Appendix D Datasets
The following is a complete description of the datasets used in this work:
  • Text2Image: A cross-modal dataset, where data and query points may have different distributions in a shared space [Simhadri et al., 2022]. We use a subset consisting of $10$ million $200$-dimensional data points along with a subset of 10,00010{,}000 test queries. The dataset is available under the terms of the Creative Commons Attribution 4.0 International license.
  • Music: $1$ million $100$-dimensional points [Morozov and Babenko, 2018] with 1,0001{,}000 queries. To the best of our knowledge, the dataset comes with no information on the license under which it is made available (per https://github.com/stanis-morozov/ip-nsw?tab=readme-ov-file).
  • DeepImage: Subset of $10$ million $96$-dimensional points from the billion deep image features dataset [Yandex and Lempitsky, 2016] with 10,00010{,}000 queries. The dataset is available under the terms of Apache license 2.0.
  • GloVe: 1.21{.}2 million, $200$-dimensional word embeddings trained on tweets [Pennington et al., 2014] with 10,00010{,}000 test queries. The dataset is available under the terms of the Public Domain Dedication and license v1.0.
  • MsMarco-MiniLM: Ms Marco Passage Retrieval v1 [Nguyen et al., 2016] is a question-answering dataset consisting of 8.88{.}8 million short passages in English. We use the “dev” set of queries for retrieval, made up of 6,9806{,}980 questions. We embed individual passages and queries using the all-MiniLM-L6-v2 model [^10] Checkpoint at https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. to form a $384$-dimensional vector collection. The dataset is available under the terms of the Creative Commons Attribution 4.0 International license. The model used to embed the dataset is available under the terms of Apache license 2.0.
  • Nq-Ada2: 2.72{.}7 million, 1,5361{,}536-dimensional embeddings of the Natural Questions dataset [Kwiatkowski et al., 2019] with the Ada-002 model. [^11] https://openai.com/index/new-and-improved-embedding-model/ The dataset is available under the terms of Apache license 2.0.
@@ Appendix D Datasets
Appendix E Effect of hyper-parameters on Optimist
@@
Figure 6: Top-$100$ recall vs. volume of probed data as we change the rank parameter ($t$). $\textsc{Optimist}(0,\cdot)$ is a sketch that is the diagonals only. Partitioning is with Spherical KMeans.
@@
Figure 7: Top-$100$ recall vs. volume of probed data, comparing a range of values of $\delta$. As $\delta\rightarrow 1$, Optimist becomes more optimistic. Partitioning is with Spherical KMeans.
@@
Appendix F Experiments with spherical KMeans
(a) Nq-Ada2
(b) Music
(c) GloVe
Figure 8: Top-$k$ recall vs. volume of probed data. Partitioning is with Spherical KMeans. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$.
(d) Text2Image
(e) MsMarco-MiniLM
(f) DeepImage
Figure 8: Top-$k$ recall vs. volume of probed data. Partitioning is with Spherical KMeans. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$.
Appendix G Experiments with standard KMeans
(a) Nq-Ada2
(b) Music
(c) GloVe
Figure 9: Top-$k$ recall vs. volume of probed data. Partitioning is with Standard KMeans. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$.
(d) Text2Image
(e) MsMarco-MiniLM
(f) DeepImage
Figure 9: Top-$k$ recall vs. volume of probed data. Partitioning is with Standard KMeans. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$.
Appendix H Experiments with GMM
(a) Music
(b) GloVe
(c) Text2Image
(d) MsMarco-MiniLM
Figure 10: Top-$k$ recall vs. volume of probed data. Partitioning is with Gaussian Mixture Model. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$. We note that, due to the dimensionality of Nq-Ada2, we were unable to complete GMM clustering on this particular dataset.
(e) DeepImage
Figure 10: Top-$k$ recall vs. volume of probed data. Partitioning is with Gaussian Mixture Model. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$. We note that, due to the dimensionality of Nq-Ada2, we were unable to complete GMM clustering on this particular dataset.
Appendix I Latency comparison
Refer to caption Refer to caption Refer to caption
(a) 11 thread
(b) 22 threads
(c) 44 threads
Figure 11: Mean latency (ms) to reach $95\%$ recall, using different number of threads (columns), when PQ-compressed shards are on SSD (top row) and blob storage (bottom row). For each dataset, we plot the latency breakdown for NormalizedMean (top bar) and Optimist (bottom bar), and report relative gains (negative value indicates gain by Optimist).
Appendix J Maximum inner product prediction
@@ Appendix I Latency comparison
Figure 12: Mean prediction error $\mathcal{E}_{\ell}(\tau,\cdot)$, defined in Equation (9), versus $\ell$ (expressed as percent of total number of shards), for various routers and datasets.